OpenStudy (anonymous):

Use the definition of derivative to find f'(x) for f(x)=1/((3x+2)^.5)

OpenStudy (amistre64):

rewrite it so that its just a power rule and a chain rule

OpenStudy (anonymous):

I can't--my work has to show the definition of a derivative, so I have to take limh-->0 (f(x+h)-f(x))/h

OpenStudy (amistre64):

its still rather simple, just long winded

OpenStudy (amistre64):

\[\Large\frac{\frac{1}{(3(x+h)+2)^.5}-\frac{1}{(3x+2)^.5}}{h}\] \[\Large\frac{\frac{1}{\sqrt{(3x+3h+2)}}-\frac{1}{\sqrt{(3x+2)}}}{h}\] \[\frac{\sqrt{(3x+2)}-\sqrt{(3x+3h+2)}}{h\sqrt{(3x+2)}\sqrt{(3x+3h+2)}}\] \[\frac{\sqrt{(3x+2)}-\sqrt{(3x+3h+2)}}{h\sqrt{(3x+2)(3x+3h+2)}}\] \[\frac{\sqrt{(3x+2)}-\sqrt{(3x+3h+2)}}{h\sqrt{(9x^2+9xh+12x+6h+4)}}\]very involved :)

OpenStudy (amistre64):

might need to use a conjugate of the top to try to pull out an h to cancel with

OpenStudy (anonymous):

Ok--I had multiplied by the conjugate right after the first step and things got messy. This is a little bit less messy.

OpenStudy (amistre64):

i really dont see the need to torture students with this messy stuff once the derivative rules have been proven by simpler problems

OpenStudy (anonymous):

my prof regularly does stuff like this. in real analysis...

OpenStudy (amistre64):

\[\frac{\sqrt{(3x+2)}-\sqrt{(3x+3h+2)}}{h\sqrt{(3x+2)}\sqrt{(3x+3h+2)}}\] \[\frac{-3h}{h\sqrt{(3x+2)}\sqrt{(3x+3h+2)}(conj)}\] \[\frac{-3}{\sqrt{(3x+2)}\sqrt{(3x+3h+2)}(conj)}\] \[\frac{-3}{\sqrt{(3x+2)}\sqrt{(3x+2)}~(\sqrt{(3x+2)}+\sqrt{(3x+2)}}\] \[\frac{-3}{(3x+2)~(2\sqrt{(3x+2)}}\] maybe?

OpenStudy (amistre64):

\[(3x+2)^{-1/2}\] \[-\frac12(3x+2)^{-3/2}*3\] \[\frac{-3}{2(3x+2)^{3/2}}\] \[\frac{-3}{2(3x+2)\sqrt{(3x+2)}}\]

OpenStudy (anonymous):

yep, that's it. Thanks!

OpenStudy (amistre64):

;) good luck